# Haifeng Ruan*       Yuntong Zhang*       Abhik Roychoudhury

### v3

@@ Abstract
Autonomous program improvement typically involves automatically producing bug fixes and feature additions. Such program improvement can be accomplished by a combination of large language model (LLM) and program analysis capabilities, in the form of an LLM agent. Since program repair or program improvement typically requires a specification of intended behavior - specification inference can be useful for producing high quality program patches. In this work, we examine efficient and low-cost workflows for iterative specification inference within an LLM agent. Given a GitHub issue to be resolved in a software project, our goal is to conduct iterative code search accompanied by specification inference - thereby inferring intent from both the project structure and behavior. The intent thus captured is examined by a reviewer agent with the goal of vetting the patches as well as providing a measure of confidence in the vetted patches. Our approach SpecRover (AutoCodeRover-v2) is built on the open-source LLM agent AutoCodeRover. In an evaluation on the full SWE-Bench consisting of 2294 GitHub issues, it shows more than 50% improvement in efficacy over AutoCodeRover. Compared to the open-source agents available, our work shows modest cost (\$0.65 per issue) in resolving an average GitHub issue in SWE-Bench lite. The production of explanation by SpecRover allows for a better “signal” to be given to the developer, on when the suggested patches can be accepted with confidence. SpecRover also seeks to demonstrate the continued importance of specification inference in automated program repair, even as program repair technologies enter the LLM era.
Index Terms:
Large Language Models, Autonomous Software Engineering, Program Repair, Specification Inference, Explainable AI in coding.
@@ I Introduction
  • Experimental evidence: Our tool shows high efficacy, solving $19.3\%$ issues in full SWE-bench and $31\%$ on SWE-bench lite. We also balance other needs from LLM agents such as low cost (\$0.65 per issue) and supporting higher precision/recall.
@@ V-C RQ3: Quality of Specifications produced
In this section, we illustrate the quality of evidence generated by SpecRover with two examples. The high-quality evidence allows a developer to easily integrate auto-generated patch into an existing codebase.
@@ VII Related Work
Automated program repair (APR) [8, 21] is a well studied research area in software engineering. Given a buggy program $P$, and a test-suite $T$, automated program repair attempts to (minimally) modify $P$ to a program $P^{\prime}$ which passes the given test-suite $T$. APR techniques involve metaheuristic search [22], semantic analysis [6], machine learning [23], or a combination of different techniques. APR can also be used to rectify automatically generated code from LLMs, e.g. see the work of [1] among others.
@@ VIII Perspectives
Owing to the growth of large language based automatic programming (see [24] for a recent summary), there exists interest in autonomous program improvement technologies. We propose SpecRover with the perspective of autonomously producing patches which are suggested with confidence (thus developers can confidently accept them) and come with explanations. The technical innovations supporting SpecRover are the specification inference to guide patching, and the rigorous vetting of patches via our reviewer agent. Our work on SpecRover seeks to put the matter of quality of patches produced by LLM agents into the research community’s attention, whereas other works are mostly focusing on the agent efficacy.
Tool Release
We will share access to the tool within 45 days of the paper first appearing in public domain via arXiv.
Appendix A Supplementary Material: Data Memorization Investigation
An issue that can bias the evaluation of LLM agent-generated patches is data memorization. Data memorization occurs when an LLM deals with a program that exists in its training set. This issue is rather common, since LLMs have been trained on a vast amount of publicly available code. In the presence of possible data memorization, when a correct patch is generated by an LLM agent, it may be not because of the capability of the LLM agent, but because the LLM has seen the repaired version of the program. In this way, the capability of the LLM agent is overestimated.
One way of gauging the threat of data memorization is to check whether the generated patch is syntactically equivalent to the ground-truth correct patch. Intuitively, the LLM is likely to generate a syntactically equivalent patch if it has seen the correct program. Therefore, the threat of data memorization is lower if the generated patch turns out not syntactically equivalent.
To gauge the data memorization threat and have a more accurate evaluation of the LLM agents, we perform this syntactic check of patches generated by agent systems on SWE-Bench lite. For each issue resolved by a tool, we compare the patch generated by the tool against the ground-truth developer-written patch, and check whether they are syntactically equivalent. The results are shown in Table III. The column “# Resolved” shows the number of resolved issues by a tool; “# Syntactically Equiv.” shows the number of resolved issues with the generated patch being syntactically equivalent to the ground-truth patch; “Equiv. / Resolved” shows the ratio between these two numbers.
Tool # Resolved # Syntactically % Equiv.
Equiv. / Resolved
SpecRover 93 9 10%
Agentless 82 14 17%
Moatless Tools 80 12 15%
Aider 79 6 8%
OpenDevin 78 0 0%
AppMap Navie 65 8 12%
AutoCodeRover 57 3 5%
SWE-agent 54 8 15%
TABLE III: Number of resolved instances, and number of resolved instances with generated patches being syntactically equivalent to the ground-truth patch. Results are shown for each tool in SWE-Bench lite.
(a) Issue statement.
(b) Ground-truth patch.
Figure 10: Issue Django-11163. The issue statement contains the ground-truth patch.
Discussion on the results
On SWE-Bench lite, SpecRover generated a correct patch for 93 issues. Out of these, only 9 patches (10%) are syntactically identical with the ground-truth patch. Most other tools show similar percentages of syntactically identical patches, ranging from 5% to 17%. OpenDevin has 0% identical patches based on our computation, because OpenDevin often creates backup files for some of the files in the codebase, and these backup changes are also contained in its patch.
Overall, the threat of data memorization is minor in our evaluation of SpecRover. We note that the 10% identical patches represent an upper bound of the effect of potential data memorization. In some issues, the issue reporter has suggested the correct patch. An example of such issues is shown in Figure 10. For these issues, it is natural that the LLM can produce an identical patch, irrespective of data memorization. Moreover, even if an identical patch is produced, it can be that the correct patch is simple, which can be easy for an LLM to produce such a patch.

### v4

@@ Abstract
Autonomous program improvement typically involves automatically producing bug fixes and feature additions. Such program improvement can be accomplished by a combination of large language model (LLM) and program analysis capabilities, in the form of an LLM agent. Since program repair or program improvement typically requires a specification of intended behavior - specification inference can be useful for producing high quality program patches. In this work, we examine efficient and low-cost workflows for iterative specification inference within an LLM agent. Given a GitHub issue to be resolved in a software project, our goal is to conduct iterative code search accompanied by specification inference - thereby inferring intent from both the project structure and behavior. The intent thus captured is examined by a reviewer agent with the goal of vetting the patches as well as providing a measure of confidence in the vetted patches. Our approach SpecRover is built on the open-source LLM agent AutoCodeRover. In an evaluation on the full SWE-Bench consisting of 2294 GitHub issues, it shows more than 50% improvement in efficacy over AutoCodeRover. Compared to the open-source agents available, our work shows modest cost (\$0.65 per issue) in resolving an average GitHub issue in SWE-Bench lite. The production of explanation by SpecRover allows for a better “signal” to be given to the developer, on when the suggested patches can be accepted with confidence. SpecRover also seeks to demonstrate the continued importance of specification inference in automated program repair, even as program repair technologies enter the LLM era.
@@ I Introduction
  • Experimental evidence: Our tool shows high efficacy, solving $19.3\%$ issues in full SWE-bench and $31\%$ on SWE-bench lite. We also balance other needs from LLM agents such as low cost (\$0.65 per issue) and supporting higher precision/recall. Our tool is available open-source in Zenodo and will be publicly released upon acceptance.
@@ IV Experimental Setup
Manual Inspection of Results
For a better understanding of the experimental results, we perform manual inspection to obtain certain data, e.g., patch overfitting rate. All manual inspections were independently conducted by two authors and subsequently cross-checked. Discrepancies were resolved through consultation with a third colleague from our group.
@@ V-A RQ1: Overall Efficacy of Task Resolving
Data Memorization
An issue that can bias the evaluation of LLM agent-generated patches is data memorization [19]. Data memorization occurs when an LLM deals with a program that exists in its training set. To evaluate the risk of data leakage, we specifically count the patches generated by SpecRover that are syntactically identical to the ground-truth patches. Our counting shows that in SWE-Bench Lite, SpecRover generated the ground-truth patch only for 9 issues, accounting for 10% of the 93 resolved issues. The result shows that data memorization occurs very infrequently for SpecRover.
@@
Patch Correctness
A patch that passes a given test suite is not necessarily correct, because the test suite is an incomplete specification. This problem is known as patch overfitting [21] in program repair. For a more accurate evaluation of the generated patches, we manually compared the SpecRover-generated patch with the developer patch for each resolved issue. Our manual investigation confirmed that 60.2% (56/93) of the patches that pass the test suites are semantically equivalent to the ground-truth patches. We also observed that 29 out of the 37 (=93-56) overfitting patches (i.e., the patches that pass the test suite but are not semantically equivalent to ground truth) modify the same methods as the ground-truth patch, and only the specific modification is semantically different from the ground truth. This implies that even overfitting patches produced by SpecRover are useful in indicating a fix location. Adding up the semantically equivalent patches and the overfitting patches that have the correct fix location, a total of over 90% of the test-passing patches prove useful for issue resolution.
Reasons for Failure
We examined the 207 tasks that are not resolved by SpecRover in SWE-Bench Lite.These failure cases break down into three cases:
  1. 1.
    Ambiguous issue description (as labeled by SWE-bench Verified [22]): 107 (51.7%)
  2. 2.
    Incorrect fix location: 61 (29.5%)
  3. 3.
    Incorrect code modification: 39 (18.8%)
Case 1 is “hard or impossible to solve” [22] without further information. Case 2 is more frequent than Case 3, implying that SpecRover can potentially have significant improvement by leveraging more fix localization techniques. For Case 3, in most such cases (26/39), we observed that the gist of the patch generated by SpecRover is actually correct, but some code detail is wrong. To reduce this kind of failures, more test generation techniques can be leveraged within our agent.
@@ V-C RQ3: Quality of Specifications produced
In this section, we investigate the quality of evidence generated by SpecRover. The high-quality evidence allows a developer to easily integrate auto-generated patch into an existing codebase.
V-C1 Function Summaries
We manually investigated the natural-language specifications generated by SpecRover for SWE-bench Lite issues. To obtain some approximation of “ground truth” for the generated specifications, we extracted the titles of Pull Requests that developers wrote when they fixed each issue. The PR title usually summarizes the fixes made in the PR in one sentence. In this investigation, we include only the issues for which SpecRover generated a patch at the same methods as the developer’s patch. Moreover, we exclude the issues whose developer-written PR title is too brief (e.g. “Fixes issue #…”). For the included issues, we manually compare the specification generated by SpecRover for the patched methods, with the PR title written by the developer. For 72/101 (71%) inspected issues, the specification generated by SpecRover covers similar intent to the human-written PR title, which shows that the generated specification serves as a good summary of the intended behavior. The generated specifications are generally an extended version of the PR title and additionally describe how the code should be modified.
V-C2 Reviewer Feedback
We cannot conduct a quantitative examination on the quality of the reviewer feedbacks on the generated patches, since there is no “ground truth”. Therefore, we illustrate the quality of the feedbacks with two examples.
@@ VII Related Work
Automated program repair (APR) [8, 24] is a well studied research area in software engineering. Given a buggy program $P$, and a test-suite $T$, automated program repair attempts to (minimally) modify $P$ to a program $P^{\prime}$ which passes the given test-suite $T$. APR techniques involve metaheuristic search [25], semantic analysis [6], machine learning [26], or a combination of different techniques. APR can also be used to rectify automatically generated code from LLMs, see e.g. [1].
@@
Previous works have studied function-level specification inference by means of LLMs [27, 28, 29]. However, these works either focus on generating specifications for simple one-function or one-class programs [28, 27], or assume that a target function is provided [29]. In contrast, SpecRover targets large programs and does not assume the target functions to generate specifications for are given.
Since both tests and patches generated in an autonomous workflow can be unreliable, we additionally consider the natural-language issue description when judging their correctness. The judgement process is concretized as the reviewer agent. Compared to the previous works that discover additional specification based on user intent [30], we do not rely on interactive user feedback. Instead, we utilize the high-level natural-language description as additional feedback. Furthermore, compared to previous works that evaluate the generated code with an LLM-based reviewer [31], our reviewer agent is designed for the setup where both natural-language instructions and unreliable tests are present.
@@ VIII Perspectives
Owing to the growth of LLM-based automatic programming (see [32] for a recent summary), there exists interest in autonomous program improvement technologies. We propose SpecRover with the perspective of autonomously producing patches which are suggested with confidence (thus developers can confidently accept them) and come with explanations. The technical innovations supporting SpecRover are the specification inference to guide patching, and the rigorous vetting of patches via our reviewer agent. Our work on SpecRover seeks to put the matter of quality of patches produced by LLM agents into the research community’s attention, whereas other works are mostly focusing on the agent efficacy.
Moving forward, we envision that LLM agents will need to improve the precision and recall. Specifically, LLM agents will need to vet the produced patches. The vetting needs to be accompanied by sophisticated test generation, so that there is a rich test-suite to check the produced patches. Thus, we need to take the viewpoint of a developer using an LLM agent, who would be concerned about (a) efficacy, (b) cost, and, most importantly (c) signal-to-noise ratio in the agent output.
Data Availability
We share full public access of the source code and experimental artifacts of SpecRover at https://zenodo.org/doi/10.5281/zenodo.13161650.
We are continuously improving the efficacy and usability of SpecRover (i.e., AutoCodeRover-v2). At the time of paper acceptance (Nov 2024), AutoCodeRover-v2 has achieved an efficacy of 37.3% on SWE-bench Lite and 46.2% on SWE-bench Verified. We have also packaged AutoCodeRover-v2 as a GitHub bot offering one-click issue resolution [33]. The latest source code, experimental results, and news updates of AutoCodeRover-v2 can be found at our GitHub repository [^5] https://github.com/nus-apr/auto-code-rover and website [^6] https://www.autocoderover.net/.
Acknowledgments
This work was partially supported by a Singapore Ministry of Education (MoE) Tier 3 grant ”Automated Program Repair”, MOE-MOET32021-0001.